Performance Gloves, Inc. produces three sizes of sports gloves: small, medium, and large. A glove pattern is first stenciled onto leather in the Pattern Department. The stenciled patterns are then sent to the Cut and Sew Department, where the glove is cut and sewed together. Performance Gloves uses the multiple production department factory overhead rate method of allocating factory overhead costs. Its factory overhead costs were budgeted as follows: Pattern Department overhead $176,800 297,000 Cut and Sew Department overhead $473,800 Total The direct labor estimated for each production department was as fol lows: Pattern Department 2,600 direct labor hours Cut and Sew Department 3,300 5,900 direct labor hours Total Direct labor hours are used to allocate the production department overhead to the products. The direct labor hours per unit for each product for each production department were obtained from the engineering records as follows: Production Departments Small Glove Medium Glove Large Glove Pattern Department 0.04 0.05 0.06 Cut and Sew Department 0.08 0.10 0.12 Direct labor hours per unit 0.12 0.15 0.18 If required, round all per unit answers to the nearest cent.
a. Determine the two production department factory overhead rates. per dlh Pattern Department per dlh Cut and Sew Department b. Use the two production department factory overhead rates to determine the factory overhead per unit for each product. Small glove per unit Medium glove per unit Large glove per unit
Answer #1

a) Overhead rate:

pattern department = 176800/2600 = 68 per DLH

Cut and Sew department = 297000/3300 = 90 per DLH

b) Factory overhead per unit

Small glove = (68*.04+90*.08) = 9.92 per unit

Medium glove = (68*.05+90*.10) = 4.30 per unit

Large glove = (68*.06+90*.12) = 14.88 per unit